# Elasto-frictional reduced model of a cyclically sheared container filled with particles

### v1

@@ Abstract
This article explores the hysteretic behavior and the damping features of sheared granular media using discrete element method (DEM) simulations. We consider polydisperse non-cohesive frictional spherical particles, enclosed in a container with rigid but moving walls, submitted to a cyclic simple shear superimposed to a confining pressure. The mechanical response of the grains is analyzed in the permanent regime, by fitting the macroscopic stress-strain relation applied to the box with a Dahl-like elasto-frictional model. The influence of several parameters such as the amplitude of the strain, the confining pressure, the elasticity, the friction coefficient, the size and the number of particles are explored. We find that the fitted parameters of our macroscopic Ansatz qualitatively rely on both a well-established effective medium theory of confined granular media and a well-documented rheology of granular flow. Quantitatively, we demonstrate that the single degree-of-freedom elasto-frictional reduced model reliably describes the nonlinear response of the granular layer over a wide range of operating conditions. In particular, we show that the mechanical response of a granular slab under simple shear depends on an unique dimensionless parameter, akin to an effective Coulomb threshold, at low shear/high pressure. Further, exploring higher shear/lower pressure, we evidence an optimal damping at the crossover between a loose unjammed regime and a dense elastic regime.
@@ 1 Introduction
Granular materials are ubiquitous in Nature, where they are found in a variety of states of matter (Jaeger et al., 1996). From the application point of view, granular media are known to dissipate and disperse efficiently mechanical shocks (a sand pile can stop a riffle bullet) and they proved to be efficient also in the context of sound and vibration mitigation (Melo et al., 2006), damping (Masmoudi et al., 2016) and insulation (Hentati et al., 2021). Compared to classical matter, granular media present unusual properties like softening (Johnson and Jia, 2005), weakening (Jia et al., 2011), phase transition (Zuñiga et al., 2022) relaxation (Ren et al., 2013) and aging (Espíndola et al., 2016), to name but a few.
@@
Jammed granular media are frequently referred to as granular solids owing to their ability to resist to pressure and shear, or even as metamaterials (Götz and Pöschel, 2023) to tailor wave absorption (Vasconcellos et al., 2022) and propagation (Vasconcellos et al., 2024). To determine their elastic features, an effective medium theory (EMT) was developed in the late 80’s to connect the elastic moduli at the particle level to the one at the sample scale (Walton, 1987). Such a macroscopic description of a pile of grains inherit and reveal the intrinsically nonlinear Hertzian potential at the heart of the contacts mechanics (Johnson, 1985; Popov, 2010). When the particles are non-cohesive and dry, they are likely submitted also to friction (Popov, 2010). Frictional properties are particularly studied in the field of granular flow, in line with many industrial applications, where one of the bottleneck relies on understanding how friction dictates the rheology of an avalanche, for instance (GDR MiDi, 2004). Here, a key description is found in dense and slow granular flows, where such a behavior can be described with the tools and definitions of soil mechanics (Rognon et al., 2008; Jop, 2015; Fall et al., 2017). In this context, an effective friction coefficient is defined as the ratio between shear stress and confining pressure (GDR MiDi, 2004). Interestingly, this macroscopic coefficient only depends on a dimensionless inertial number $I$ (Da Cruz et al., 2005; Jop et al., 2006).
@@
Inter-particle friction, together with inelastic collisions between grains, is also a key feature to damp mechanical energy in discrete systems. Indeed, vibrated granular media are able to achieve perfectly inelastic collapse conditions (McNamara and Young, 1992; Pacheco-Vázquez and Dorbolo, 2013), leading to exceptional damping efficiency (Sánchez et al., 2012; Michon et al., 2013; Masmoudi et al., 2016; Ferreyra et al., 2021). Here, the dry friction being mostly a frequency independent mechanism, it is much more effective at low frequency than the viscoelastic damping (Kausel and Assimaki, 2002; Rouleau, 2013). Phenomenologically, the friction between numerous particles gives rise to powerful hysteretic behaviors at the sample scale (Bi et al., 2011; Ren et al., 2013), which themselves rely on multiple stick-slip events at the contact scale, leading in practice to nonlinear damping properties (Assimaki et al., 2000; Brandes, 2010; Senetakis et al., 2013; Awdi et al., 2023). The macroscopic hysteretic response is thus closely related to the anisotropy and the topological disorder of the contact network, as evidenced for instance by X-Ray tomography of granular media under shear (Xing et al., 2021).
@@
In the vast majority of the studies on granular media, simulations based on discrete element method (DEM) are widely used to model soils (Khalili et al., 2017a, b) and structure made of particles (Sitharam and Vinod, 2010; Zhong and Sun, 2018) to explore in-depth phenomena at the microscopic scale; such a detailed description generally remains a challenging task in experiments, due to the large number of interacting bodies. For instance, Luding et al (Magnanimo and Luding, 2011; Krijgsman and Luding, 2013) used DEM simulations in order to build a constitutive model describing the behavior of a bi-axial box filled with particles submitted to pure shear in two-dimensions (Magnanimo and Luding, 2011), by taking into account for various phenomena like anisotropy, hysteresis and ratcheting. This model was later enriched to include pressure relaxation and stabilization in pure shear (Krijgsman and Luding, 2013), allowing to describe transient as well as permanent regimes. Other extensive models for granular matter include granular solid hydrodynamics (Jiang and Liu, 2009) handling static behavior with plasticity, but also transient loading (Gudehus, 2019) including hysteresis.
@@
From the practical engineering point of view, models were also derived to describe the macroscopic mechanical response of structural elements, such as a slender composite beam containing a granular medium core, from either simulations (Götz and Pöschel, 2023) or experiments (Bajkowski et al., 2019). In line with these attempts, our study aims at deriving a simplified but reliable description of a vibrated confined granular medium, capable of reducing the complexity of DEM simulations in order to facilitate, for instance, the optimization of vibrations dampers containing particles (Fichant et al., 2024). In particular, we require our reduced model to be parametrized by identifiable and measurable quantities, such as the features of the grains (size, material, …) and the features of the mechanical driving (confinement pressure, shear amplitude, …). Starting from a DEM numerical model of individual grains confined in a box and submitted to cyclic simple shear, we thus provide and analyze, in this study, an elasto-hysteretic reduced model at the macroscopic scale. The evolution of the elastic as well as the damping features of both the numerical model and the reduced model are explored, seeking in particular for configurations with high damping ratio.
@@
The paper is organized as follows. The Sec. 2 describes the setup of the DEM simulation to model a pack of grains enclosed in a box undergoing simple shear cyclically. In Sec. 3, the evolution of quantities of interest such as the shear stress, the pressure, the elasticity and frictional behaviors are identified and examined at the light of an elasto-hysteretic Ansatz made of a hysteretic term rendered by Dahl model, combined to a linear elastic response. the reliability of this model is probed via an extensive parametric study. In Sec. 4, we evidence how our Ansatz relates to an effective medium theory describing the macroscopic elasticity of granular media, and to a rheological model describing granular flows. Further, we find through a dimensional analysis that the granular material can be described reasonably by a single non-dimensional parameter, akin to an effective Coulomb threshold. The reliability of this model is discussed and then used to estimate the damping features. Finally, we propose a nonlinear improvement of the model, stemming from the anisotropy-like pressure fluctuations observed at high shear/low pressure; these fluctuations are described by a second non-dimensional parameter. The Sec. 5 is the conclusion.
@@ 2 Method and protocol
We build a numerical version of a granular medium with a lateral size that is large enough compared to the average particle radius, in order to get rid of the discreetness of the medium at the scale of the representative elementary volume. Our numerical sample thus consists of $1000$ ($10\times 10\times 10$) polydisperse and purely elastic spherical particles (nominally made of steel: mass density $\rho_{p}=7850$ kg/m3, Young’s modulus $E_{p}=210$ GPa, Poisson ratio $\nu_{p}=0.27$, inter-particle friction coefficient $\mu_{p}=0.15$), with uniformly distributed radii to prevent crystallization ($r_{p}=1\pm 0.25$ mm). Grains are placed within a cubic box with rigid but movable walls. The model is simulated using DEM with Yade (S̆milauer and et al., 2021) open source software. For this, Newton’s equations of motion are integrated in the time domain with an explicit second order Velocity Verlet scheme, with a constant time step chosen equal to the time of flight of the longitudinal waves across the smallest particle, $\Delta t=\mathrm{min}(r_{p}/\sqrt{M_{p}/\rho_{p}})$, being $M_{p}=E_{p}(1-\nu_{p})/(1+\nu_{p})(1-2\nu_{p})$ the P-wave modulus of particles material. This time step is always smaller than the duration of inter-particle mechanical contact (i.e. the smallest duration in the system) because the deformation rate is smaller than the P-wave speed (Johnson, 1985; Popov, 2010); this choice thus ensures a proper discretization in time to fulfill energy conservation. Elasto-frictional interactions between particles are described by Hertz-Mindlin contact law coupled with a Mohr-Coulomb criterion, which means that the normal and the tangential contact forces $F_{n,t}$ versus deformations $\delta_{n,t}$ take the form (Johnson, 1985; Popov, 2010)
@@
if the particles overlap due to compression and stick due to friction; otherwise, both forces tend to zero when the non-cohesive particles do not overlap, $\delta_{n}\leq 0$, and the tangential force saturates if sliding occurs, $F_{t}=\mathrm{min}(\|F_{t}\|,\mu_{p}F_{n})\times\mathrm{sign}(\dot{\delta}_{t})$. Here, $G_{p}=E_{p}/2(1+\nu_{p})$ is the shear modulus of the particles material and $k_{n,t}$ are their normal and tangential contact stiffnesses. Without loss of generality, the particle-to-wall interactions are chosen frictionless ($\mu_{w}=0$). This choice implies that no shear stress occurs between the particles and the walls in any directions, and simplifies the interpretation on the origin of the dissipation which, in this case, only comes from inter-particle friction. This approximation is supported by the fact that the particle-to-wall contacts are far less numerous than the particle-to-particle ones, they introduce a negligible correction in the macroscopic mechanical response.
@@
In a first step, we prepare a granular sample. We generate the $1000$ size-distributed particles placed at random positions in a dilute state (at rest and well separated), inside a large cubic box delimited by walls, using an Yade’s builtin generator. Then, we set the sample under pressure, by proceeding with an isotropic compression until a given volume fraction $\phi=V_{p}/V$ is reached, with $V_{p}$ the material volume of the spheres and $V$ the volume of the parallelepiped box. This step is achieved by imposing the displacement of all the walls at a constant and slow normal velocity, see Fig. 1 (top-left); note the walls do not interact with each other.
@@
In a second step, the granular packing is submitted to cyclic simple shear. In practice, the box is deformed by imposing the rotation of two of the side walls, see Fig. 1 (top-right), similarly to what was done experimentally in (Xing et al., 2021). However, in our case, the volume of the box is kept constant, whereas the confining pressure is kept constant in (Xing et al., 2021). We impose an harmonic rotation angle $\theta(t)=\theta_{m}\sin(\omega_{0}t)$ leading to a shear strain $\gamma=\tan(\theta)$ which is quasi-harmonic in the limit of small angles, $\gamma(t)\simeq\gamma_{m}\sin{(\omega_{0}t)}$ with $\gamma_{m}\simeq\theta_{m}\ll 1$. It is important to note that, here and in the following, cyclic shear transformations are quasi-static in order to neglect inertia in the forces and moments balance. The latter condition is $\omega_{0}t_{w}\ll 1$ in practice, the angular frequency being proportional to the shear rate, $\omega_{0}\propto\dot{\gamma}$, and the time of flight of a compressional wave across the granular medium being $t_{w}=r_{p}/\sqrt{M_{\mathrm{eff}}/\rho_{\mathrm{eff}}}$, with $\rho_{\mathrm{eff}}\propto\rho_{p}$ the mass density and $M_{\mathrm{eff}}=K_{\mathrm{eff}}+(4/3)G_{\mathrm{eff}}$ the longitudinal modulus of the sample’s effective medium defined in Sec. 4.1 below.
@@
Figure 1: (top) Granular sample during (left) initial compression and (right) cyclic shear. (a) Imposed shear strain, (b) shear stress and (c) hydrostatic pressure, as functions of time for two values of the volume fraction $\phi=0.64$ (red) and $\phi=0.70$ (blue) at $\gamma_{m}=0.15$. Stress-strain curves for (d) $\phi=0.70$ and (e) $\phi=0.64$ at various amplitude: $\gamma_{m}=0.05$, $0.10$ and $0.15$.
[…]

### v2

@@ Abstract
This article explores the hysteretic behavior and the damping features of sheared granular media using discrete element method (DEM) simulations. We consider polydisperse non-cohesive frictional spherical particles, enclosed in a container with rigid but moving walls, subjected to a cyclic simple shear superimposed on a confining pressure. The mechanical response of the grains is analyzed in the permanent regime, by fitting the macroscopic stress-strain relation applied to the box with a Dahl-like elasto-frictional model. The influence of several parameters such as the amplitude of the strain, the confining pressure, the elasticity, the friction coefficient, the size and the number of particles are explored. We find that the fitted parameters of our macroscopic Ansatz rely qualitatively on both a well-established effective medium theory of confined granular media and a well-documented rheology of granular flow. Quantitatively, we demonstrate that the single degree-of-freedom elasto-frictional reduced model reliably describes the nonlinear response of the granular layer over a wide range of operating conditions. In particular, we show that the mechanical response of a granular slab under simple shear depends on a unique dimensionless parameter akin to an effective Coulomb threshold at low shear/high pressure. Furthermore, exploring higher shear/lower pressure, we evidence optimal damping at the crossover between a loose unjammed regime and a dense elastic regime.
@@ 1 Introduction
Granular materials are ubiquitous in Nature, where they are found in a variety of states of matter (Jaeger et al., 1996). From the application point of view, granular media are known to dissipate and disperse mechanical shocks efficiently (a sand pile can stop a rifle bullet) and they have also proved to be efficient in the context of sound and vibration mitigation (Melo et al., 2006), damping (Masmoudi et al., 2016) and insulation (Hentati et al., 2021). Compared to classical matter, granular media present unusual properties like softening (Johnson and Jia, 2005), weakening (Jia et al., 2011), phase transition (Zuñiga et al., 2022), relaxation (Ren et al., 2013) and aging (Espíndola et al., 2016), to name but a few.
@@
Jammed granular media are frequently referred to as granular solids owing to their ability to resist pressure and shear, or even as metamaterials (Götz and Pöschel, 2023) to tailor wave absorption (Vasconcellos et al., 2022) and propagation (Vasconcellos et al., 2024). To determine their elastic features, an effective medium theory (EMT) was developed in the late 1980s to connect elastic moduli at the particle level to that at the sample scale (Walton, 1987). Such a macroscopic description of a pile of grains acquires and reveals the intrinsically nonlinear Hertzian potential at the heart of contact mechanics (Johnson, 1985; Popov, 2010). When the particles are non-cohesive and dry, they are also likely subjected to friction (Popov, 2010). Frictional properties are studied in particular in the field of granular flow, in line with many industrial applications, where one of the bottlenecks relies on understanding how friction dictates the rheology of an avalanche, for instance (GDR MiDi, 2004). Here, a key description is found in dense and slow granular flows, where such behavior can be described with the tools and definitions of soil mechanics (Rognon et al., 2008; Jop, 2015; Fall et al., 2017). In this context, an effective friction coefficient is defined as the ratio between shear stress and confining pressure (GDR MiDi, 2004). Interestingly, this macroscopic coefficient depends only on a dimensionless inertial number $I$ (Da Cruz et al., 2005; Jop et al., 2006).
@@
Inter-particle friction, together with inelastic collisions between grains, is also a key feature to damp mechanical energy in discrete systems. Indeed, vibrated granular media are able to achieve perfectly inelastic collapse conditions (McNamara and Young, 1992; Pacheco-Vázquez and Dorbolo, 2013), leading to exceptional damping efficiency (Sánchez et al., 2012; Michon et al., 2013; Masmoudi et al., 2016; Ferreyra et al., 2021). Here, as dry friction is mostly a frequency independent mechanism, it is much more effective at low frequency than viscoelastic damping (Kausel and Assimaki, 2002; Rouleau, 2013). Phenomenologically, the friction between numerous particles gives rise to powerful hysteretic behaviors at the sample scale (Bi et al., 2011; Ren et al., 2013), which themselves rely on multiple stick-slip events at the contact scale, leading in practice to nonlinear damping properties (Assimaki et al., 2000; Brandes, 2010; Senetakis et al., 2013; Awdi et al., 2023). The macroscopic hysteretic response is thus closely related to the anisotropy and the topological disorder of the contact network, as evidenced for instance by X-Ray tomography of granular media under shear (Xing et al., 2021).
@@
In the vast majority of the studies on granular media, simulations based on the discrete element method (DEM) are widely used to model soils (Khalili et al., 2017a, b) and structures made of particles (Sitharam and Vinod, 2010; Zhong and Sun, 2018) to explore in-depth phenomena at the microscopic scale; such a detailed description generally remains a challenging task in experiments, due to the large number of interacting bodies. For instance, Luding et al (Magnanimo and Luding, 2011; Krijgsman and Luding, 2013) used DEM simulations to build a constitutive model describing the behavior of a bi-axial box filled with particles subjected to pure shear in two-dimensions (Magnanimo and Luding, 2011), by taking into account various phenomena like anisotropy, hysteresis and ratcheting. This model was later enriched to include pressure relaxation and stabilization in pure shear (Krijgsman and Luding, 2013), making it possible to describe transient as well as permanent regimes. Other extensive models for granular matter include granular solid hydrodynamics (Jiang and Liu, 2009) handling static behavior with plasticity, as well as transient loading (Gudehus, 2019) including hysteresis.
@@
From the practical engineering point of view, models were also derived to describe the macroscopic mechanical response of structural elements, such as a slender composite beam containing a granular medium core, from either simulations (Götz and Pöschel, 2023) or experiments (Bajkowski et al., 2019). In line with these attempts, our study aims at deriving a simplified but reliable description of a vibrated confined granular medium, capable of reducing the complexity of DEM simulations to facilitate, for instance, the optimization of vibration dampers containing particles (Fichant et al., 2024). In particular, we require our reduced model to be parametrized by identifiable and measurable quantities, such as the features of the grains (size, material, etc.) and the features of mechanical action (confinement pressure, shear amplitude, etc.). Starting from a DEM numerical model of individual grains confined in a box and subjected to cyclic simple shear, in this study we thus provide and analyze a reduced elasto-hysteretic model at the macroscopic scale. The evolution of the elastic and the damping features of both the numerical model and the reduced model are explored, seeking in particular for configurations with a high damping ratio.
@@
The paper is organized as follows. Sec. 2 describes the setup of the DEM simulation to model a pack of grains enclosed in a box undergoing simple shear cyclically. In Sec. 3, the evolution of quantities of interest such as shear stress, pressure, elasticity and frictional behaviors are identified and examined at the light of an elasto-hysteretic Ansatz composed of a hysteretic term rendered by a Dahl model combined with a linear elastic response. The reliability of this model is probed via an extensive parametric study. In Sec. 4, we evidence how our Ansatz relates to an effective medium theory describing the macroscopic elasticity of granular media, and to a rheological model describing granular flows. Furthermore, we find through a dimensional analysis that the granular material can be described reasonably by a single non-dimensional parameter, akin to an effective Coulomb threshold. The reliability of this model is discussed and then used to estimate the damping features. Finally, we propose a nonlinear improvement of the model, stemming from the anisotropy-like pressure fluctuations observed at high shear/low pressure; these fluctuations are described by a second non-dimensional parameter. Sec. 5 provides the conclusion.
@@ 2 Method and protocol
We build a numerical version of a granular medium with a lateral size that is large enough compared to the average particle radius, to eliminate the discreetness of the medium at the scale of the representative elementary volume. Considerations on the number, the size and the polydispersity of the particles are detailed in the Appendix Sec. A4. Our numerical sample thus consists of $1000$ ($10\times 10\times 10$) polydisperse and purely elastic spherical particles (nominally made of steel: mass density $\rho_{p}=7850$ kg/m3, Young’s modulus $E_{p}=210$ GPa, Poisson ratio $\nu_{p}=0.27$, inter-particle friction coefficient $\mu_{p}=0.15$), with uniformly distributed radii to prevent crystallization ($r_{p}=1\pm 0.25$ mm). The grains are placed within a cubic box with rigid but movable walls. The model is simulated using DEM with Yade (S̆milauer and et al., 2021) open source software. To do this, Newton’s equations of motion are integrated in the time domain with an explicit second order Velocity Verlet scheme, with a constant time step chosen equal to the time of flight of the longitudinal waves across the smallest particle, $\Delta t=\mathrm{min}(r_{p}/\sqrt{M_{p}/\rho_{p}})$, being $M_{p}=E_{p}(1-\nu_{p})/(1+\nu_{p})(1-2\nu_{p})$ the P-wave modulus of particles material. This time step is always smaller than the duration of inter-particle mechanical contact (i.e. the smallest duration in the system) because the deformation rate is smaller than the P-wave speed (Johnson, 1985; Popov, 2010); this choice thus ensures a proper discretization in time to fulfill energy conservation. Elasto-frictional interactions between particles are described by Hertz-Mindlin contact law coupled with a Mohr-Coulomb criterion, which means that the normal and the tangential contact forces $F_{n,t}$ versus deformations $\delta_{n,t}$ take the form (Johnson, 1985; Popov, 2010)
@@
if the particles overlap due to compression and stick due to friction; otherwise, both forces tend to zero when the non-cohesive particles do not overlap, $\delta_{n}\leq 0$, and the tangential force saturates if sliding occurs, $F_{t}=\mathrm{min}(\|F_{t}\|,\mu_{p}F_{n})\times\mathrm{sign}(\dot{\delta}_{t})$. Here, $G_{p}=E_{p}/2(1+\nu_{p})$ is the shear modulus of the particles’ material and $k_{n,t}$ are their normal and tangential contact stiffnesses. Without loss of generality, the particle-to-wall interactions are chosen as frictionless ($\mu_{w}=0$). This choice implies that no shear stress occurs between the particles and the walls in any directions, and simplifies the interpretation on the origin of the dissipation which, in this case, only comes from inter-particle friction.
@@
In the first step, we prepare a granular sample. We generate the $1000$ size-distributed particles placed at random positions in a dilute state (at rest and well separated), inside a large cubic box delimited by walls, using a Yade builtin generator. Then, we set the sample under pressure, by proceeding with an isotropic compression until a given volume fraction $\phi=V_{p}/V$ is reached, with $V_{p}$ being the material volume of the spheres and $V$ the volume of the parallelepiped box. This step is achieved by imposing the displacement of all the walls at a constant and slow normal velocity, see Fig. 1 (top-left); note the walls do not interact with each other.
@@
In the second step, the granular packing is subjected to cyclic simple shear. In practice, the box is deformed by imposing the rotation of two of the side walls, see the $yz$ planes in Fig. 1 (top-right), similarly to what was done experimentally in (Xing et al., 2021). However, in our case, the volume of the box is kept constant, whereas the confining pressure is kept constant in (Xing et al., 2021). Thanks to the choice of a frictionless particle-to-wall interaction, the numerical model can be simplified by defining the four other walls, parallel to $xy$ or $xz$ planes in Fig. 1, as rigid and motionless. Otherwise, accounting for the parietal friction would require considering a deformable container to accommodate realistically the relative motion between grains and walls, which is beyond the aim of the present study and a simple DEM formulation. We impose a harmonic rotation angle (in radians) $\theta(t)=\theta_{m}\sin(\omega_{0}t)$ of the two $yz$ walls around the $z$-axis leading to a shear strain $\gamma=\tan(\theta)$ which is quasi-harmonic in the limit of small angles, $\gamma(t)\simeq\gamma_{m}\sin{(\omega_{0}t)}$ with $\gamma_{m}\simeq\theta_{m}\ll 1$ rad. It is important to note that, here and in the following, cyclic shear transformations are quasi-static in order to neglect inertia in the balance of forces and moments. The latter condition is $\omega_{0}t_{w}\ll 1$ in practice, as the angular frequency is proportional to the shear rate, $\omega_{0}\propto\dot{\gamma}$, and the time of flight of a compressional wave across the granular medium is $t_{w}=r_{p}/\sqrt{M_{\mathrm{eff}}/\rho_{\mathrm{eff}}}$, with $\rho_{\mathrm{eff}}\propto\rho_{p}$ the mass density and $M_{\mathrm{eff}}=K_{\mathrm{eff}}+(4/3)G_{\mathrm{eff}}$ the longitudinal modulus of the sample’s effective medium defined in Sec. 4.1 below.
@@
Figure 1: (top) Granular sample during (left) initial compression and (right) cyclic shear. (a) Imposed shear strain, (b) shear stress and (c) hydrostatic pressure, as functions of time for two values of the volume fraction $\phi=0.64$ (red) and $\phi=0.70$ (blue) at $\gamma_{m}=0.15$. Stress-strain curves for (d) $\phi=0.70$ and (e) $\phi=0.64$ at various amplitudes: $\gamma_{m}=0.05$, $0.10$ and $0.15$.
The DEM computation allows extracting the instantaneous quantities at the contact scale from which we compute the variables of interests and the state variables as functions of time: namely, the shear strain $\gamma(t)$, the shear stress $\tau(t)$, the hydrostatic pressure $P(t)$, the average coordination number $Z(t)$, and the volume fraction $\phi(t)=\mathrm{const}$ (maintained constant during the simple shear transformation). The pressure in the granular medium is estimated as the trace of the stress tensor $P=(\sigma_{xx}+\sigma_{yy}+\sigma_{zz})/3$, where $\sigma_{ii}=(\mathbf{F}_{+i}\cdot\mathbf{n}_{+i}+\mathbf{F}_{-i}\cdot\mathbf{n}_{-i})/2S_{i}$ with $S_{i}$ being the area of face $i=(x,y,z)$. It is computed from the total forces $\mathbf{F}_{\pm i}$ exerted by the particles on the faces defined by their normal vectors $\mathbf{n}_{\pm i}$. We checked that this estimation of the confining pressure is consistent, by less than 0.10.1% relative difference, with that computed from the trace of the Love-Weber stress tensor (Makse et al., 2004; Thornton, 2000; S̆milauer and et al., 2021). The onset of jamming (O’Hern et al., 2003) (which corresponds to the onset of a pressure in the sample, at which it gets an elasticity) is reached at around a volume fraction $\phi_{c}=0.595\pm 0.003~(0.5\%)$ on average (Faulconnier, 2023, p.111), over the entire dataset presented in this study. The shear stress $\tau$ is computed by summing all the reactions applied by the particles on the side walls, which are equivalent to two resisting moments computed at the center of rotation of the walls. The moments $\mathbf{M}_{+x}^{Oz}$ and $\mathbf{M}_{-x}^{Oz}$, exerted on the two opposite faces $\mathbf{n}_{\pm x}$, are used to compute the magnitude of an equivalent shear stress $\tau=-2(\mathbf{M}_{+x}^{Oz}\cdot\mathbf{z}+\mathbf{M}_{-x}^{Oz}\cdot\mathbf{z})/V$} exerted in-plane on the top wall $\mathbf{n}_{+z}$, where $\mathbf{z}$ is the Cartesian vector. Here, as a crude check, we verify that the work done by the shear force (the injected energy) equals the total energy per unit volume in the granular medium computed by the DEM, $\int{\tau d\gamma}=e$.